# The Sunburst Arc with JWST: III. An Abundance of Direct Chemical Abundances

### v2

@@ Abstract
We measure the gas-phase abundances of the elements He, N, O, Ne, S, Ar, and Fe in the Lyman-continuum emitting region of the Sunburst Arc, a highly magnified galaxy at redshift . We detect the temperature-sensitive auroral lines [Sii], [Oiii], [Siii], [Oii], and [Neiii] in a stacked spectrum of 5 multiple images of the Lyman-continuum emitter (LCE), from which we directly measure the electron temperature in the low, intermediate, and high ionization zones. We also detect the density-sensitive doublets of [Oii], [Sii], and [Ariv], which constrain the density in both the low- and high-ionization gas. With these temperature and density measurements, we measure gas-phase abundances with similar rigor as studies of local galaxies. We measure a gas-phase metallicity for the LCE of , and find an enhanced nitrogen abundance . This nitrogen abundance is consistent with enrichment from a population of Wolf-Rayet stars, additional signatures of which are reported in a companion paper. Abundances of sulfur, argon, neon, and iron are consistent with local low-metallicity HII regions and low-redshift galaxies. This study represents the most complete chemical abundance analysis of a galaxy at Cosmic Noon to date, which enables direct comparisons between local HII regions and those in the distant universe.
@@ 1 Introduction
As stars shed mass in their late evolutionary stages and explode as supernovae, they recycle newly synthesized elements into the surrounding gas, enriching the nebula. Analysis of the chemical abundance patterns in nebulae thus give insight into the evolution of stars and the star formation histories of galaxies.
Recent measurements of high redshift galaxies have found unusual abundance patterns, including anomalously high nitrogen abundances (N/O, Cameron et al., 2023; Senchyna et al., 2023; Marques-Chaves et al., 2024; Castellano et al., 2024) and carbon abundances (C/O, D’Eugenio et al., 2023). Meanwhile similar carbon abundance measurements at slightly lower redshifts () find much lower abundances (Citro et al., 2023; Jones et al., 2023). The anomalous abundance patterns in the highest redshift galaxies may be indicative of chemical evolution via e.g. Wolf-Rayet (WR) stars or extremely metal poor stars, as the standard production pathway for nitrogen and carbon via intermediate-mass AGB stars would not have time to develop.
One of the most robust methods to determine chemical abundance patterns in galaxies is the so-called “direct” method, which directly measures electron temperature () and density from the strengths of auroral emission lines relative to bright collisionally-excited lines. These temperatures and densities are then used in conjunction with the strengths of the bright lines to calculate the abundance of each ion within the nebula (Dinerstein, 1990).
The auroral lines are intrinsically faint, making this measurement difficult. Direct abundances have been well studied in nearby galaxies (e.g., van Zee & Haynes, 2006; Berg et al., 2015; Croxall et al., 2015, 2016; Berg et al., 2020; Rogers et al., 2021, 2022). In these local galaxies, multiple auroral lines are detected, allowing direct measurements of the electron temperature in multiple ionization zones within the nebular gas. This eliminates the need for empirical temperature relations to infer the physical conditions of different zones within the gas, thereby making the final abundance determinations more accurate. Additionally, including muliple ionization zones minimizes the impact of empirically-calibrated ionization corrections, which account for unobserved ionization states of a given element (e.g., Izotov et al., 2006; Amayo et al., 2021). In the best case, all relevant ionization states of each element are observed, resulting in the most accurate determinations of elemental abundances.
@@
This paper presents the analysis of direct abundances in the Sunburst Arc (Dahle et al., 2016), a highly magnified gravitationally lensed galaxy at redshift (Sharon et al., 2022; Diego et al., 2022; Pignataro et al., 2021). We focus on a single physical region of the galaxy that is leaking ionizing photons (Rivera-Thorsen et al., 2019), which we refer to as the Lyman-continuum emitter (LCE). This region is a compact star cluster (Vanzella et al., 2022) hosting a young stellar population with a light-weighted age of Myr (Chisholm et al., 2019; Rivera-Thorsen et al., 2024) and a steep UV slope (Kim et al., 2023). Several authors have found evidence of very massive stars in the LCE (Chisholm et al., 2019; Meštrić et al., 2023; Pascale et al., 2023). A companion paper presents evidence that the LCE hosts a population of Wolf-Rayet stars (Rivera-Thorsen et al., 2024).
In this paper, we report the detection of the [Sii], [Oiii], [Siii], [Oii], and [Neiii] auroral lines, from which we measure directly in the low, intermediate, and high ionization zones of the nebula. We also measure the electron temperature in ionized hydrogen gas from the Balmer jump. We measure abundances for seven elements (He, N, O, Ne, S, Ar, Fe). For three of these (O, S, Ar) we measure multiple ionization states, which minimizes the impact of ionization corrections in our abundance determinations. With this set of diagnostics, for the first time we apply the same level of rigor used in local galaxies to an object at Cosmic Noon, bridging the gap between local abundance measurements and those at high redshift.
@@ 2 Data
The Sunburst Arc was observed by JWST in Cycle 1 using NIRCam (Rieke et al., 2023) and the Integral Field Spectroscopy (IFS) mode of NIRSpec (Böker et al., 2023), taken in program GO-02555 (PI: Rivera-Thorsen). NIRCam data were taken in six broadband filters (F115W, F150W, F200W, F277W, F356W, and F444W). NIRSpec data were taken with two grating settings (G140H/F100LP, G235H/F170LP). The IFS data were taken in three pointings, shown in Figure 1 of Rivera-Thorsen et al. (2024). The observations and data reduction are described in greater detail in Rivera-Thorsen et al., in prep; we briefly summarize the relevant steps here. The NIRSpec IFS data were reduced following the methods described in Rigby et al. (2023b), using the TEMPLATES NIRSpec data reduction notebook (Rigby et al., 2024). We used the JWST data reduction pipeline version 1.11.4 (Bushouse et al., 2023), along with calibration reference files from pmap_1105. After running the main JWST pipeline, we use the baryon-sweep111https://github.com/aibhleog/baryon-sweep code to remove any remaining outliers from the final data cubes (Hutchison et al., 2023, 2024).
In this work, we analyze a spatially-integrated spectrum that consists of 5 multiple images of the LCE covered by the three IFU pointings. We extract spectra of the LCE images from spaxel apertures, using a flux-weighted sum. The spectra are normalized, no magnification correction is applied, however all results presented herein rely exclusively on flux ratios which are unaffected by magnification or normalization. The full extraction method is documented in greater detail in Rivera-Thorsen et al. (2024).
@@ 3.2 Emission Line Measurements
Emission lines are fit using Gaussian profiles, utilizing the publicly-available tools developed by the TEMPLATES team and available on GitHub.222github.com/JWST-Templates We initially attempt to fit each line with a single Gaussian component, and iteratively add up to two additional components if visual inspection determines the fit to be poor. We find that the brightest lines (e.g. [Oiii], H) require three components - one narrow (unresolved) component, one medium component with a full-width at half maximum (FWHM) of km s-1, and one broad component with a FWHM of km s-1. As line strengths decrease, the broadest component is the first to become lost in the noise. Thus for many lines, only the unresolved and the medium width components are used in fitting. For the faintest lines (including all of the auroral lines), only a single unresolved component is sufficient to obtain a good fit. When fitting both single- and multi-component lines, we allow the centroids of each component to vary within Å in the observed frame to allow for possible velocity offsets from our assumed systemic redshift ( Rivera-Thorsen et al., 2024). For bright multi-component lines, the amplitude and width of the Gaussians are left unconstrained. For the faintest emission features, we choose to fix the line widths to the instrumental resolution at the observed wavelength, as is observed in the narrow components of the brighter emission lines. This prevents noise around the edge of the lines from artificially broadening the Gaussian fits. Final line fluxes are obtained by integrating the Gaussian fits. For line fits with two components, we sum the Gaussian integrals to obtain the total line flux. For three component fits, we choose to not include the broadest component in our total flux measurements, as the velocities ( km s-1) are such that this gas has likely traveled far from the main nebula. This only affects the H and [Oii] lines. Uncertainties in measured line fluxes are propagated through from fitting uncertainties on the Gaussian amplitude and width.
@@ 3.3 Extinction Correction
Emission line fluxes are first corrected for Milky Way dust reddening using the dust law of Cardelli et al. (1989). We query the Milky Way dust map of Schlafly & Finkbeiner (2011) via the NASA/IPAC Infrared Science Archive333https://irsa.ipac.caltech.edu/applications/DUST/ to obtain the ISM reddening correction. We find the Galactic .
We next correct for dust within the LCE itself. This correction is done using an iterative process using the H, H, H, and H line strengths. We start by calculating the expected Balmer line ratios, assuming an initial electron temperature of K, and an initial electron density of cm-3. The measured line ratios are then compared to the predicted values to calculate , assuming the dust law of Cardelli et al. (1989). We then apply the error-weighted average dust correction, weighted by the uncertainty on the Balmer line ratios, to relevant emission lines, and calculate the electron temperatures using the [Oiii], [Oii], and [Siii] lines. The electron density is calculated using the [Sii] doublet ratio, as that is the best-measured of the density-sensitive doublets used in this study. We update the electron temperature using an error-weighted average of the three calculated temperatures and repeat the dust correction measurement. This process is continued iteratively until the change in error-weighted electron temperature is less than 15 K. The reddening that we calculated from this iterative process is . The reddening-corrected line intensities are reported in Table 1.
@@ 4.1 Electron Temperature and Density
The broad wavelength coverage (rest-frame Å) and high SNR of the stacked LCE spectrum enables calculations of direct electron temperatures and densities from multiple ions. We detect the auroral emission lines of [Oiii], [Oii], [Siii], and [Sii], which are shown in Figure 1. These let us use the temperature-sensitive auroral-to-nebular line ratios to calculate the electron temperatures for O++, O+, S++, and S+. Our spectra do not go red enough to cover the [Siii] line, so the S++ temperature is measured from [Siii]. While we detect both [Sii] lines, the redder line is barely detected (). We therefore choose to use the ratio [Sii] to calculate the temperature for this ion. We do not detect the [Nii] line, and instead set an upper limit on its strength. This sets an upper limit on the ratio [Nii], and thus an upper limit on the N+ temperature. We also detect the density-sensitive lines [Sii], [Oii], and [Ariv]. We calculate the electron temperatures and densities from these lines using the getTemDen function from the python package PYNEB (Luridiana et al., 2015). The resulting measurements are presented in Table 3.
@@ 4.1.1 Balmer Jump Temperature
Our spectrum covers the Balmer jump at Å, and the high SNR of the stacked spectrum allow us to detect this continuum feature. The strength of the jump is sensitive to the electron temperature of the ionized hydrogen in the nebula. We measure the jump following the procedure of Liu et al. (2001). The continuum redward and blueward of the jump is measured with a linear fit (shown in Figure 2). On the red side of the jump, we mask emission lines prior to fitting. We also mask the region between the jump at Å and the H12 line, where blended high-order Balmer lines will bias the continuum level. We evaluate the blue linear fit at 3643Å, and the red linear fit at 3681Å.
Following Equation 3 of Liu et al. (2001), the Balmer jump temperature depends on the ratio of the jump strength to the H11 line flux, and the helium abundances He+/H+ and He++/H+. Our helium abundance measurements are described in greater detail in the following section. We do not detect nebular Heii; while the line is present, we only see evidence of a broad component originating in the WR stars (see Rivera-Thorsen et al., 2024). We therefore assume the He++/H. The dependence on helium abundance is weak (see also Fig. 9 of Liu et al., 2001). As a test of this weak dependence, if we assume the He++ abundance is 50% of the He+, the calculated temperature does not change beyond our quoted uncertainties.
With this method, we measure a Balmer jump temperature of K. This temperature is largely consistent with our other measured electron temperatures.

### v3

@@ Abstract
We measure the gas-phase abundances of the elements He, N, O, Ne, S, Ar, and Fe in an individual Hii region known to be leaking Lyman-continuum photons in the Sunburst Arc, a highly magnified galaxy at redshift . We detect the temperature-sensitive auroral lines [Sii], [Oiii], [Siii], [Oii], and [Neiii] in a stacked spectrum of 5 multiple images of the Lyman-continuum emitter (LCE), from which we directly measure the electron temperature in the low, intermediate, and high ionization zones. We also detect the density-sensitive doublets of [Oii], [Sii], and [Ariv], which constrain the density in both the low- and high-ionization gas. With these temperature and density measurements, we measure gas-phase abundances with similar rigor as studies of local galaxies and Hii regions. We measure a gas-phase metallicity for the LCE of , and find an enhanced nitrogen abundance . This nitrogen abundance is consistent with enrichment from a population of Wolf-Rayet stars, additional signatures of which are reported in a companion paper. Abundances of sulfur, argon, neon, and iron are consistent with local low-metallicity Hii regions and low-redshift galaxies. This study represents the most complete chemical abundance analysis of an individual Hii region at Cosmic Noon to date, which enables direct comparisons between local Hii regions and those in the distant universe.
@@ 1 Introduction
As stars shed mass in their late evolutionary stages and explode as supernovae, they recycle newly synthesized elements into the surrounding gas, enriching the nebula. Analysis of the chemical abundance patterns in nebulae thus gives insight into the evolution of stars and the star formation histories of galaxies.
Recent measurements of high redshift () galaxies have found unusual abundance patterns, including anomalously high nitrogen abundances (N/O, Cameron et al., 2023; Senchyna et al., 2023; Marques-Chaves et al., 2024; Castellano et al., 2024) and carbon abundances (C/O, D’Eugenio et al., 2023). Meanwhile, similar carbon abundance measurements at slightly lower redshifts () find much lower abundances (Citro et al., 2023; Jones et al., 2023). The anomalous abundance patterns in the highest redshift galaxies may be indicative of chemical evolution via e.g. Wolf-Rayet (WR) stars or extremely metal poor stars, as the standard production pathway for nitrogen and carbon via intermediate-mass AGB stars would not have time to develop.
One of the most robust methods to determine chemical abundance patterns in galaxies is the so-called “direct” method, which directly measures electron temperature () and density from the strengths of auroral emission lines relative to bright collisionally excited lines. These temperatures and densities are then used in conjunction with the strengths of the bright lines to calculate the abundance of each ion within the nebula (Dinerstein, 1990).
The auroral lines are intrinsically faint, making this measurement difficult. Direct abundances have been well studied in nearby galaxies (e.g., van Zee & Haynes, 2006; Berg et al., 2015; Croxall et al., 2015, 2016; Berg et al., 2020; Rogers et al., 2021, 2022). In these local galaxies, multiple auroral lines are detected, allowing direct measurements of the electron temperature in multiple ionization zones within the nebular gas. This eliminates the need for empirical temperature relations to infer the physical conditions of different zones within the gas, thereby making the final abundance determinations more accurate. Additionally, including multiple ionization zones minimizes the impact of empirically calibrated ionization corrections, which account for unobserved ionization states of a given element (e.g., Izotov et al., 2006; Amayo et al., 2021). In the best case, all relevant ionization states of each element are observed, resulting in the most accurate determinations of elemental abundances.
@@
Measurements of direct abundances in distant galaxies have thus far only been made using the integrated light of an entire galaxy. Studies of nearby galaxies have the advantage of resolving individual Hii regions within galaxies and measuring the abundances for these regions independently, finding that abundances tend to vary across galactocentric radius (e.g., Zaritsky et al., 1994; Moustakas et al., 2010; Sánchez-Menguiano et al., 2018; Berg et al., 2020). Abundance measurements using galaxy-integrated spectra therefore do not provide a complete picture of the abundance structure within the galaxy.
Abundance patterns of sub-galactic objects can also shed light on their growth and development. For example, multiple stellar populations found in globular clusters show clear abundance trends between the first and second generations of stars (e.g., Bastian & Lardo, 2018). The exact mechanism of enrichment for second generation stars is debated, with fast rotating massive stars, AGB stars, binary interactions, and supermassive stars (among others) proposed as the sources of enrichment (see e.g. Charbonnel (2016) for a review). Measurements of individual Hii regions and proto-globular cluster candidates in distant galaxies can help to shed additional light on the questions of chemical evolution at the sub-galactic scale and the growth of globular clusters.
This paper presents the analysis of direct abundances in the Sunburst Arc (Dahle et al., 2016), a highly magnified gravitationally lensed galaxy at redshift (Sharon et al., 2022; Diego et al., 2022; Pignataro et al., 2021). We focus on a single physical region of the galaxy that is leaking ionizing photons (Rivera-Thorsen et al., 2019), which we refer to as the Lyman-continuum emitter (LCE). This region is a compact star cluster (Vanzella et al., 2022) hosting a young stellar population with a light-weighted age of Myr (Chisholm et al., 2019; Rivera-Thorsen et al., 2024) and a steep UV slope (Kim et al., 2023). Several authors have found evidence of very massive stars in the LCE (Chisholm et al., 2019; Meštrić et al., 2023; Pascale et al., 2023). A companion paper presents evidence that the LCE hosts a population of Wolf-Rayet (WR) stars (Rivera-Thorsen et al., 2024).
In this paper, we report the detection of the [Sii], [Oiii], [Siii], [Oii], and [Neiii] auroral lines, from which we measure directly in the low, intermediate, and high ionization zones of the nebula. We also measure the electron temperature in ionized hydrogen gas from the Balmer jump. We measure abundances for seven elements (He, N, O, Ne, S, Ar, Fe). For three of these (O, S, Ar) we measure multiple ionization states, which minimizes the impact of ionization corrections in our abundance determinations. With this set of diagnostics, for the first time we apply the same level of rigor used in local Hii regions to an object of similar physical scale at Cosmic Noon, bridging the gap between local abundance measurements and those at high redshift.
@@ 2 Data
The Sunburst Arc was observed by JWST in Cycle 1 using NIRCam (Rieke et al., 2023) and the Integral Field Spectroscopy (IFS) mode of NIRSpec (Böker et al., 2023), taken in program GO-02555 (PI: Rivera-Thorsen).
NIRCam data were taken in six broadband filters (F115W, F150W, F200W, F277W, F356W, and F444W).
NIRSpec data were taken with two grating settings (G140H/F100LP, G235H/F170LP).
The IFS data were taken in three pointings, shown in Figure 1 of Rivera-Thorsen et al. (2024).
The observations and data reduction are described in greater detail in Rivera-Thorsen et al., in prep; we briefly summarize the relevant steps here.
The NIRSpec IFS data were reduced following the methods described in Rigby et al. (2023b), using the TEMPLATES NIRSpec data reduction notebook (Rigby et al., 2024).
We used the JWST data reduction pipeline version 1.11.4 (Bushouse et al., 2023), along with calibration reference files from pmap_1105.
After running the main JWST pipeline, we use the baryon-sweep111https://github.com/aibhleog/baryon-sweep,
DOI: 10.5281/zenodo.8377531 code to remove any remaining outliers from the final data cubes (Hutchison et al., 2023, 2024).
In this work, we analyze a spatially-integrated spectrum that consists of 5 multiple images of the LCE covered by the three IFU pointings. We extract spectra of the LCE images from spaxel apertures, using a flux-weighted sum. The spectra are normalized by the median flux in the wavelength region overlapping between the two gratings. No magnification correction is applied, however all results presented herein rely exclusively on flux ratios which are unaffected by magnification or normalization. Normalized spectra from the individual apertures are summed to create the final stacked spectrum. The full extraction method is documented in greater detail in Rivera-Thorsen et al. (2024).
@@ 3.2 Emission Line Measurements
Emission lines are fit using Gaussian profiles, utilizing the publicly-available tools developed by the TEMPLATES team and available on GitHub.222github.com/JWST-Templates We initially attempt to fit each line with a single Gaussian component, and iteratively add up to two additional components if visual inspection determines the fit to be poor. We find that the brightest lines (e.g. [Oiii], H) require three components - one narrow (unresolved) component, one medium component with a full-width at half maximum (FWHM) of km s-1, and one broad component with a FWHM of km s-1. The width of the medium component is consistent with measurements from Mainali et al. (2022), and all components are consistent with Rivera-Thorsen et al. (2024). As line strengths decrease, the broadest component is the first to become lost in the noise. Thus for many lines, only the unresolved and the medium width components are used in fitting. For the faintest lines (including all of the auroral lines), only a single unresolved component is sufficient to obtain a good fit. When fitting both single- and multi-component lines, we allow the centroids of each component to vary within Å in the observed frame to allow for possible velocity offsets from our assumed systemic redshift ( Rivera-Thorsen et al., 2024). For bright multi-component lines, the amplitude and width of the Gaussians are left unconstrained. For the faintest emission features, we choose to fix the line widths to the instrumental resolution reported in the JDox dispersion files333https://jwst-docs.stsci.edu/jwst-near-infrared-spectrograph/nirspec-instrumentation/nirspec-dispersers-and-filters at the observed wavelength, as is observed in the narrow components of the brighter emission lines. This prevents noise around the edge of the lines from artificially broadening the Gaussian fits. However we note that the NIRSpec IFU dispersions have yet to calibrated based on in-flight measurements, so it is possible that this method introduces a small systematic uncertainty. Final line fluxes are obtained by integrating the Gaussian fits. For line fits with two components, we sum the Gaussian integrals to obtain the total line flux. For three component fits, we choose to not include the broadest component in our total flux measurements, as the velocities ( km s-1) are such that this gas has likely traveled far from the main nebula. This only affects the H and [Oiii] lines. Uncertainties in measured line fluxes are propagated through from fitting uncertainties on the Gaussian amplitude and width.
@@
The [Ariv] line is blended with a neighboring Hei emission line at Å. To correct for the contribution from Hei, we calculate the expected flux of this line based on the nearby Hei line, assuming a temperature of 15000 K and a density of 1200 cm-2. The calculated strength of this Hei line is reported in Table 1, with uncertainty propagated from the Hei flux measurement. The calculated Hei is subtracted from the original measured [Ariv] line flux, and the uncertainties are added in quadrature to produce the final [Ariv] flux reported in Table 1.
The [Nii] line is not detected in our spectrum. We set a upper limit on this line flux using the summed flux in a neighboring region with no emission or absorption lines present. We randomly perturb the flux density in each wavelength step in this region by sampling a Gaussian distribution centered on the measured flux density, with the standard deviation set by the flux density uncertainty. We then sum the randomly perturbed flux in the region. We repeat this process 1000 times, then use the 99.85th percentile of the resulting summed flux distribution as our upper limit. The resulting upper limit is reported in Table 1.
We test our measured upper limit in two ways. First, we inject mock emission lines into the blank region used for the [Nii] upper limit calculation, then attempt to fit these injected lines with a Gaussian profile. We find that a 3 measurement from the mock line matches our upper limit calculation. As a second consistency check, we calculate the upper limit using blank spectral regions near our other two faintest lines, [Neiii] and [Sii]. We find that the upper limits for these lines are consistent with the measured fluxes for these lines. From these tests, we conclude that our upper limit is reliable.
@@ 3.3 Extinction Correction
Emission line fluxes are first corrected for Milky Way dust reddening using the dust law of Cardelli et al. (1989). We query the Milky Way dust map of Schlafly & Finkbeiner (2011) via the NASA/IPAC Infrared Science Archive444https://irsa.ipac.caltech.edu/applications/DUST/ to obtain the ISM reddening correction. We find the Galactic .
We next correct for dust within the LCE itself. This correction is done using an iterative process using the H, H, H, and H line strengths. We start by calculating the expected Balmer line ratios, assuming an initial electron temperature of K, and an initial electron density of cm-3. The measured line ratios are then compared to the predicted values to calculate , assuming the dust law of Cardelli et al. (1989). We then apply the error-weighted average dust correction, weighted by the uncertainty on the Balmer line ratios, to relevant emission lines, and calculate the electron temperatures using the [Oiii], [Oii], and [Siii] lines. The electron density is calculated using the [Sii] doublet ratio, as that is the best-measured of the density-sensitive doublets used in this study. We update the electron temperature using an error-weighted average of the three calculated temperatures and repeat the dust correction measurement. This process is continued iteratively until the change in error-weighted electron temperature is less than 15 K. The reddening that we calculated from this iterative process is . The reddening-corrected line intensities are reported in Table 1. Our measured reddening is consistent with that reported in Rivera-Thorsen et al. (2024) using the attenuation law of Calzetti et al. (2000), and with that reported from the modeling of Pascale et al. (2023). The reddening reported in Mainali et al. (2022) is somewhat higher, however skylines impacting their H measurement could be the cause of the discrepancy.
@@ 4.1 Electron Temperature and Density
The broad wavelength coverage (rest-frame Å) and high SNR of the stacked LCE spectrum enables calculations of direct electron temperatures and densities from multiple ions. We detect the auroral emission lines of [Oiii], [Oii], [Siii], [Sii], and [Neiii]which are shown in Figure 1. These let us use the temperature-sensitive auroral-to-nebular line ratios to calculate the electron temperatures for O++, O+, S++, S+, and Ne++. Our spectra do not go red enough to cover the [Siii] line, so the S++ temperature is measured from [Siii]. While we detect both [Sii] lines, the redder line is barely detected (). We therefore choose to use the ratio [Sii] to calculate the temperature for this ion. The [Neiii] line is weakly detected ( significance), leading to a larger calculated uncertainty on the Ne++ temperature. We do not detect the [Nii] line, and instead set a upper limit on its strength. This sets a corresponding upper limit on the ratio [Nii], and thus an upper limit on the N+ temperature. We also detect the density-sensitive lines [Sii], [Oii], and [Ariv]. We calculate the electron temperatures and densities from these lines using the getTemDen function from the python package PYNEB (Luridiana et al., 2015). The resulting measurements are presented in Table 3.
@@ 4.1.1 Balmer Jump Temperature
Our spectrum covers the Balmer jump at Å, and the high SNR of the stacked spectrum allow us to detect this continuum feature. The strength of the jump is sensitive to the electron temperature of the ionized hydrogen in the nebula. The continuum around the Balmer jump feature is a combination of stellar and nebular emission, so we model these two components simultaneously following the procedure of Hayes et al. (2024). The stellar component is based on Starburst99 template spectra (Leitherer et al., 1999; Vázquez & Leitherer, 2005), assuming a Salpeter (1955) IMF and a constant star formation history. The nebular spectra are created using the Schirmer (2016) code, and include contributions from recombination lines, free-bound, free-free, and 2-photon continuua. Both the stellar and nebular light are attenuated using the dust law of Calzetti et al. (2000), though we find our results do not change significantly when using alternate dust attenuation prescriptions.
When fitting, we mask out metal emission lines, as these are not included in the model. We also mask the WR bumps reported in Rivera-Thorsen et al. (2024), as the Starburst99 stellar templates do not include WR emission. The region near the NIRSpec detector gap is also masked. Finally, the spectral templates used only extend to 7000Å rest-frame, so we do not include the longer wavelengths in our fit. The data used, along with the best-fit stellar and nebular models, are shown in Figure 2.
This method yields a Balmer jump temperature of K, consistent with our other measured low-ionization temperature ( K) within . As a consistency check for our model, we compare the age of the stellar population in our fit to the age measured in Chisholm et al. (2019). Our model finds a best-fit age of Myr, consistent with the Myr measured using a similar Starburst99 model in Chisholm et al. (2019).
We previously measured the Balmer jump following the procedure of Liu et al. (2001), which provides a simple relation between Balmer jump strength and temperature assuming all the continuum is nebular in origin. With this assumption, we found a Balmer jump temperature of K. This temperature is higher than our model-based estimate, however the large uncertainty means that the two measurements are consistent within .